# The Uniform Mordell–Lang Conjecture

### v3

@@ 4.1. A result in family
Let $c_{\mathrm{deg},1}^{\prime}=\min_{1\leq i\leq k}\{c_{i,1}^{\prime}\}>0$, $c(r,d,S_{\mathrm{deg},2}^{\prime}=\max_{1\leq i\leq k}\{c_{i,2}^{\prime}\}>0$ and $c_{\mathrm{deg},3}^{\prime}=\max_{1\leq i\leq k}\{c_{i,3}^{\prime}\}>0$. We have seen above that the set $\{S_{1},\ldots,S_{k}\}$ is uniquely determined by $S$. So the constants $c_{\mathrm{deg},1}^{\prime}$, $c_{\mathrm{deg},2}^{\prime}$ and $c_{\mathrm{deg},3}^{\prime}$ thus obtained depend only on $g$, $d$, $r$ and $S$. Moreover, for each $s\in\bigcup_{i=1}^{k}S_{i}(\overline{\mathbb{Q}})$, the set
@@
must be contained in $\Sigma_{i,s}$ for some $i$. So $\Sigma_{\mathrm{deg},s}$ is contained in $X^{\prime}(\overline{\mathbb{Q}})$ for some proper Zariski closed $X^{\prime}\subsetneq\mathscr{X}_{s}$ with $\deg_{{\mathcal{L}}_{s}}(X^{\prime})\leq c_{\mathrm{deg},2}^{\prime}$. This concludes for Case (i).
Case (ii) $s\in\pi^{[m]}(\mathscr{X}_{S^{\prime}}^{[m],*})(\overline{\mathbb{Q}})$ for some irreducible component $S^{\prime}$ of $\overline{S_{\mathrm{gen}}}$. In this case, $(\mathscr{X}_{S^{\prime}}^{[m],*})_{s}\not=\emptyset$. So $\mathscr{X}_{s}^{m}=(\mathscr{X}_{S^{\prime}}^{[m]})_{s}\not=(\mathscr{X}_{S^{\prime}}^{[m]}\setminus\mathscr{X}_{S^{\prime}}^{[m],*})_{s}$.
By the height inequality of Dimitrov–Gao–Habegger [DGH21, Thm.1.6] (here, we take the version of [Gao21, Thm.7.1]), there exist constants $c=c(S^{\prime})>0$ and $c^{\prime}=c^{\prime}(S^{\prime})$ such that
@@
for all $(x_{1},\ldots,x_{m})\in(\mathscr{X}_{S^{\prime}}^{[m],*})_{s}(\overline{\mathbb{Q}})$.
Set $c_{S^{\prime},1}^{\prime}=c/m$ and $c_{S^{\prime},3}^{\prime}=(c+c^{\prime})/m$, with $c$ and $c^{\prime}$ from (4.6). Then $c_{S^{\prime},1}^{\prime}$ and $c_{S^{\prime},3}^{\prime}$ depend only on $S^{\prime}$ and $m$. Consider
@@
We claim that $\Sigma_{S^{\prime},s}^{m}\cap\mathscr{X}_{S^{\prime}}^{[m],*}(\overline{\mathbb{Q}})=\emptyset$. Indeed assume not, then there exists $(x_{1},\ldots,x_{m})\in\Sigma_{S^{\prime},s}^{m}\cap\mathscr{X}_{S^{\prime}}^{[m],*}(\overline{\mathbb{Q}})$. This contradicts the height inequality (4.6) above. So $\Sigma_{S^{\prime},s}^{m}\subseteq(\mathscr{X}_{S^{\prime}}^{[m]}\setminus\mathscr{X}_{S^{\prime}}^{[m],*})_{s}(\overline{\mathbb{Q}})$.
Recall the assumption $\mathscr{X}_{s}^{m}\not=(\mathscr{X}_{S^{\prime}}^{[m]}\setminus\mathscr{X}_{S^{\prime}}^{[m],*})_{s}$ for this case. Thus we are allowed to apply Lemma 4.3 to $X=\mathscr{X}_{s}$, $L={\mathcal{L}}_{s}|_{\mathscr{X}_{s}}$, $Z=(\mathscr{X}_{S^{\prime}}^{[m]}\setminus\mathscr{X}_{S^{\prime}}^{[m],*})_{s}$ and $\Sigma=\Sigma_{S^{\prime},s}$. So there exists a Zariski closed subset $X^{\prime}$ of $\mathscr{X}_{s}$ such that
@@
    $X^{\prime}\subsetneq\mathscr{X}_{s}$;
@@
    $\deg_{{\mathcal{L}}_{s}}(X^{\prime})<c(m,r,d,\deg_{{\mathcal{L}}_{s}^{\boxtimes m}}(\mathscr{X}_{S^{\prime}}^{[m]}\setminus\mathscr{X}_{S^{\prime}}^{[m],*})_{s})$;
@@
But $\deg_{{\mathcal{L}}_{s}^{\boxtimes m}}(\mathscr{X}_{S^{\prime}}^{[m]}\setminus\mathscr{X}_{S^{\prime}}^{[m],*})_{s}$ is bounded above solely in terms of $S^{\prime}$ and $m$. Hence property (ii) above can be simplified to be $\deg_{{\mathcal{L}}_{s}}(X^{\prime})<c_{S^{\prime},2}^{\prime}$, with $c_{S^{\prime},2}^{\prime}$ depending only on $m$, $r$, $d$ and $S^{\prime}$. In summary, the set $\Sigma_{S^{\prime},s}$ defined in (4.7) is contained in $X^{\prime}(\overline{\mathbb{Q}})$ for some proper Zariski closed $X^{\prime}\subsetneq\mathscr{X}_{s}$ with $\deg_{{\mathcal{L}}_{s}}(X^{\prime})<c_{S^{\prime},2}^{\prime}$.
Let $c_{1}^{*\prime}=\min_{S^{\prime}}\{c_{S^{\prime},1}^{\prime}\}>0$, $c_{2}^{*\prime}=\max_{S^{\prime}}\{c_{S^{\prime},2}^{\prime}\}>0$ and $c_{3}^{*\prime}=\max_{S^{\prime}}\{c_{S^{\prime},3}^{\prime}\}>0$, where $S^{\prime}$ runs over all irreducible components of $\overline{S_{\mathrm{gen}}}$; see (3.5) for the definition of $S_{\mathrm{gen}}$. The Zariski closed subset $\overline{S_{\mathrm{gen}}}$ of $S$ is uniquely determined by $S$, and hence the set $\{S^{\prime}\}$ is uniquely determined by $S$. Moreover $m=m(S)>0$ depends only on $S$. So the constants $c_{1}^{*\prime}$, $c_{2}^{*\prime}$ and $c_{3}^{*\prime}$ thus obtained depend only on $r$, $d$ and $S$. The discussion above yields the following assertion: for each $s\in\bigcup_{S^{\prime}}\pi^{[m]}(\mathscr{X}_{S^{\prime}}^{[m],*})(\overline{\mathbb{Q}})$, the set
@@
must be contained in $\Sigma_{S^{\prime},s}$ for some $S^{\prime}$; so $\Sigma_{s}^{*}$ is contained in $X^{\prime}(\overline{\mathbb{Q}})$ for some proper Zariski closed $X^{\prime}\subsetneq\mathscr{X}_{s}$ with $\deg_{{\mathcal{L}}_{s}}(X^{\prime})<c_{2}^{*\prime}$. This concludes for Case (ii).
Now let $s\in S_{\mathrm{gen}}$ be an arbitrary point. Then we are either in Case (i) or Case (ii). So the proposition follows by letting $c_{1}^{\prime}=\min\{c_{\mathrm{deg},1}^{\prime},c_{1}^{*\prime}\}>0$, $c_{2}^{\prime}=\max\{c_{\mathrm{deg},2}^{\prime},c_{2}^{*\prime}\}>0$ and $c_{3}^{\prime}=\max\{c_{\mathrm{deg},3}^{\prime},c_{3}^{*\prime}\}>0$; all the constants involved depend only on $g$, $d$ and $S$ (see (4.5) and (4.8)). ∎
The following lemma replaces [DGH21, Lem.6.3] when we work in higher dimensional cases. Let $k$ be an algebraically closed field and all varieties are assumed to be defined over $k$. Let $M\geq 1$ be an integer.
@@
Let $X$ be an irreducible projective variety with a very ample line bundle $L$. Let $Z\subsetneq X^{M}$ be a proper closed subvariety. There exists
@@
such that for any subset $\Sigma\subseteq X(k)$ satisfying $\Sigma^{M}\subseteq Z(k)$, there exists a proper closed subvariety $X^{\prime}$ of $X$ with $\Sigma\subseteq X^{\prime}(k)$ and $\deg_{L}(X^{\prime})<c$.
@@
We prove this lemma by induction on $M$. The base step $M=1$ is trivial by taking $X^{\prime}:=Z$.
Assume the lemma is proved for $1,\ldots,M-1\geq 1$. Let $q\colon X^{M}\rightarrow X$ be the projection to the first factor. Let $Y$ be any irreducible component of $Z$. Let $Z^{\prime}$ be the union of such $Y$’s with $q(Y)=X$, and $Z^{\prime\prime}$ be the union of the other components.
Set $\Sigma^{\prime\prime}:=q\left(\Sigma^{M}\cap Z^{\prime\prime}(k)\right).$ As $\Sigma^{M}\subseteq Z(k)$, we have $\Sigma^{M}\subseteq Z^{\prime}(k)\bigcup Z^{\prime\prime}(k)$. Then clearly $(\Sigma\setminus\Sigma^{\prime\prime})\times\Sigma^{M-1}\subseteq Z^{\prime}(k)$.
By the Fiber Dimension Theorem, for a generic point $P\in X(k)$, the fiber $q|_{Z^{\prime}}^{-1}(P)$ has dimension $\leq\dim Z^{\prime}-\dim X<\dim X^{M}-\dim X=\dim X^{M-1}$. Let
$W:=\{P\in X(k):\{P\}\times X^{M-1}\subseteq Z^{\prime}\}.$
Then the upper semicontinuity theorem says that $W$ is Zariski closed in $X$, which must furthermore be proper.
@@
But $\deg_{L^{\boxtimes(M-1)}}(Z_{1})=\deg_{L^{\boxtimes M}}(Z^{\prime}\cap(\{P\}\times X^{M-1}))$ and
by Bézout’s Theorem. Using (2.6) inductively by taking $Y_{1}:=X$ and $Y_{2}:=X^{i}$ for $i=1,2,...,M-2$, we have
$\deg_{L^{\boxtimes(M-1)}}(X^{M-1})=\frac{((M-1)\dim X)!}{((\dim X)!)^{M-1}}(\deg_{L}X)^{M-1}$
Moreover $\deg_{L^{\boxtimes M}}(Z^{\prime})\leq\deg_{L^{\boxtimes M}}(Z)$. Overall, we get
Hence the right hand side of (4.9) is only related to $M,\dim X,\deg_{L}X,\deg_{L^{\boxtimes M}}(Z)$.
@@
$\deg_{L}X^{\prime\prime}\leq\deg_{L^{\boxtimes M}}(X\times\{\underline{x_{0}}\})\deg_{L^{\boxtimes M}}(Z)=\deg_{L}X\deg_{L^{\boxtimes M}}(Z)$
by Bézout’s Theorem.
For $\Sigma^{\prime\prime}\subseteq q(Z^{\prime\prime})$, note that $q(Z^{\prime\prime})$ is a proper closed subset of $X$ by definition. Moreover we have
$\deg_{L}q(Z^{\prime\prime})\leq\deg_{L^{\boxtimes M}}(Z^{\prime\prime})\leq\deg_{L^{\boxtimes M}}(Z)$
Thus for this case, it suffices to take $X^{\prime}:=X^{\prime\prime}\bigcup q(Z^{\prime\prime})$. We are done. ∎
4.2. Proposition 4.1 for principally polarized abelian varieties
Proposition 4.4.
Let $g$, $r$ and $d$ be positive integers. There exist constants $c_{1}^{\prime}=c_{1}^{\prime}(g,r,d)>0$, $c_{2}^{\prime}=c_{2}^{\prime}(g,r,d)>0$ and $c_{3}^{\prime}=c_{3}^{\prime}(g,r,d)>0$ satisfying the following property. For
  • each principally polarized abelian variety $(A,L)$ defined over $\overline{\mathbb{Q}}$ of dimension $g$,
  • and each irreducible subvariety $X$ of $A$, defined over $\overline{\mathbb{Q}}$, with $\dim X=r$ and $\deg_{L}X=d$ such that $X$ generates $A$,
the set
$\Sigma:=\left\{x\in X^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{L\otimes L_{-}}(x)\leq c_{1}^{\prime}\max\{1,h_{\mathrm{Fal}}(A)\}-c_{3}^{\prime}\right\},$
where $L_{-}:=[-1]^{*}L$, satisfies the following property: $\Sigma\subseteq X^{\prime}$ for some proper Zariski closed $X^{\prime}\subsetneq X$ with $\deg_{L}(X^{\prime})<c_{2}^{\prime}$.
Proof.
Let $\mathfrak{A}_{g}\rightarrow\mathbb{A}_{g}$ be the universal abelian variety over the moduli space of principally polarized abelian varieties of dimension $g$ with level-$4$-structure.
Each principally polarized abelian variety $(A,L)$ defined over $\overline{\mathbb{Q}}$ gives rise to a point $b\in\mathbb{A}_{g}(\overline{\mathbb{Q}})$. Thus $A\cong(\mathfrak{A}_{g})_{b}$. In the rest of the proof, we will identify $A$ with $(\mathfrak{A}_{g})_{b}$.
[…]
Let $X$ be an irreducible subvariety of $A$ which generates $A$. Assume $d=\deg_{L}X$ and $r=\dim X$. We may assume $r\geq 1$ because otherwise the result is trivial. We have
Consider the restricted Hilbert scheme $\mathbf{H}^{\circ}:=\mathbf{H}^{\circ}_{r,8^{r}d}(\mathfrak{A}_{g}/\mathbb{A}_{g})$ from (3.6) and the commutative diagram (4.2)
${\mathscr{X}_{\mathbf{H}^{\circ}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{A}_{\mathbf{H}^{\circ}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$$\scriptstyle{\iota}$${\mathfrak{A}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi^{\mathrm{univ}}}$${\mathbf{H}^{\circ}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\iota_{\mathbf{H}^{\circ}}}$${\mathbb{A}_{g}}$
@@
where $\pi|_{\mathscr{X}_{\mathbf{H}^{\circ}}}\colon\mathscr{X}_{\mathbf{H}^{\circ}}\rightarrow\mathbf{H}^{\circ}$ is the universal family. All objects in this diagram are varieties defined over $\overline{\mathbb{Q}}$. Then the pair $(X,(A,L))$ is parametrized by a point $s\in\mathbf{H}^{\circ}(\overline{\mathbb{Q}})$. Thus $X=\mathscr{X}_{s}$ and $A=\mathcal{A}_{s}=(\mathfrak{A}_{g})_{b}$, and $\iota_{\mathbf{H}^{\circ}}(s)=b$.
@@
For the line bundle ${\mathcal{M}}=\overline{\iota_{\mathbf{H}^{\circ}}}^{*}\mathfrak{M}$ on $\overline{\mathbf{H}^{\circ}}$ defined above Proposition 4.2, we have $h_{\overline{\mathbf{H}^{\circ}},{\mathcal{M}}}(s)=h_{\overline{\mathbb{A}_{g}},\mathfrak{M}}(\iota_{\mathbf{H}^{\circ}}(s))=h_{\overline{\mathbb{A}_{g}},\mathfrak{M}}(b)$. By fundamental work of Faltings [MB85, Thm.1.1], $h_{\mathrm{Fal}}(A)$ is bounded from above in terms of $h_{\mathbb{A}_{g},\mathfrak{M}}(b)=h_{\overline{\mathbf{H}^{\circ}},{\mathcal{M}}}(s)$ and $g$ only. More precisely,
(4.11) $\max\{1,h_{\mathrm{Fal}}(A))\}\leq c(g)h_{\overline{\mathbf{H}^{\circ}},{\mathcal{M}}}(s)+c^{\prime}(g)\log(h_{\overline{\mathbf{H}^{\circ}},{\mathcal{M}}}(s)+2)\leq(c(g)+c^{\prime}(g))h_{\overline{\mathbf{H}^{\circ}},{\mathcal{M}}}(s)+2c^{\prime}(g)$
for $c(g)>0$ and $c^{\prime}(g)>0$.
@@
For the line bundle ${\mathcal{L}}=\iota^{*}\mathfrak{L}_{g}^{\otimes 4}$ on $\mathcal{A}_{\mathbf{H}^{\circ}}$ defined above Proposition 4.2, we have ${\mathcal{L}}|_{\mathcal{A}_{s}}=\mathfrak{L}_{g}^{\otimes 4}|_{(\mathfrak{A}_{g})_{\iota(s)}}=\mathfrak{L}_{g}^{\otimes 4}|_{(\mathfrak{A}_{g})_{b}}$, and hence ${\mathcal{L}}|_{A}=\mathfrak{L}_{g}^{\otimes 4}|_{A}$ for the identification $A=\mathcal{A}_{s}=(\mathfrak{A}_{g})_{b}$ fixed above. Thus $c_{1}({\mathcal{L}}|_{A})=4c_{1}(L\otimes L_{-})$. Moreover, both ${\mathcal{L}}|_{A}$ and $(L\otimes L_{-})^{\otimes 4}$ are symmetric and ample on $A$, so Lemma 2.3 implies

### v4

@@ 4.1. A preliminary result for restricted Hilbert schemes
Let $c_{\mathrm{deg},1}^{\prime}=\min_{1\leq i\leq k}\{c_{i,1}^{\prime}\}>0$, $c_{\mathrm{deg},2}^{\prime}=\max_{1\leq i\leq k}\{c_{i,2}^{\prime}\}>0$ and $c_{\mathrm{deg},3}^{\prime}=\max_{1\leq i\leq k}\{c_{i,3}^{\prime}\}>0$. As the set $\{S_{1},\ldots,S_{k}\}$ is uniquely determined by $S$, the constants $c_{\mathrm{deg},1}^{\prime}$, $c_{\mathrm{deg},2}^{\prime}$ and $c_{\mathrm{deg},3}^{\prime}$ depend only on $g$, $d$, $r$ and $S$. Moreover, for each $s\in\bigcup_{i=1}^{k}S_{i}(\overline{\mathbb{Q}})$, the set
@@
must be contained in $\Sigma_{i,s}$ for some $i$. So $\Sigma_{\mathrm{deg},s}$ is contained in $X^{\prime}(\overline{\mathbb{Q}})$ for some proper Zariski closed $X^{\prime}\subsetneq\mathscr{X}_{s}$ with $\deg_{{\mathcal{L}}_{s}}(X^{\prime})\leq c_{\mathrm{deg},2}^{\prime}$. This concludes the induction step in Case (i).
Case (ii): $s\in\pi^{[m]}(\mathscr{X}_{S^{\prime}}^{[m],*})(\overline{\mathbb{Q}})$ for some irreducible component $S^{\prime}$ of $\overline{S_{\mathrm{gen}}}$.
By the height inequality [DGH21, Thm. 1.6] (in the version of [Gao21, Thm. 7.1]), there exist constants $c=c(S^{\prime})>0$ and $c^{\prime}=c^{\prime}(S^{\prime})$ such that
@@
for all $(x_{1},\ldots,x_{m})\in(\mathscr{X}_{S^{\prime}}^{[m],*})_{s}(\overline{\mathbb{Q}})$. We set $c_{S^{\prime},1}^{\prime}=c/m$ and $c_{S^{\prime},3}^{\prime}=(c+c^{\prime})/m$; these constants depend only on $S^{\prime}$ and $m$. Consider
@@
If $\Sigma_{S^{\prime},s}^{m}\cap\mathscr{X}_{S^{\prime}}^{[m],*}(\overline{\mathbb{Q}})$ were non-empty, then there would exist some
$(x_{1},\ldots,x_{m})\in\Sigma_{S^{\prime},s}^{m}\cap\mathscr{X}_{S^{\prime}}^{[m],*}(\overline{\mathbb{Q}}),$
in contradiction of the height inequality (4.6). Hence $\Sigma_{S^{\prime},s}^{m}\subseteq(\mathscr{X}_{S^{\prime}}^{[m]}\setminus\mathscr{X}_{S^{\prime}}^{[m],*})_{s}(\overline{\mathbb{Q}})$. Our case assumption implies $(\mathscr{X}_{S^{\prime}}^{[m],*})_{s}\not$$\emptyset$ and thus
$\mathscr{X}_{s}^{m}=(\mathscr{X}_{S^{\prime}}^{[m]})_{s}\not=(\mathscr{X}_{S^{\prime}}^{[m]}\setminus\mathscr{X}_{S^{\prime}}^{[m],*})_{s}.$
We may hence apply Lemma 4.3 below to $X=\mathscr{X}_{s}$, $L={\mathcal{L}}_{s}|_{\mathscr{X}_{s}}$, $Z=(\mathscr{X}_{S^{\prime}}^{[m]}\setminus\mathscr{X}_{S^{\prime}}^{[m],*})_{s}$ and $\Sigma=\Sigma_{S^{\prime},s}$. It yields a Zariski closed subset $X^{\prime}$ of $\mathscr{X}_{s}$ such that
@@
    $X^{\prime}\subsetneq\mathscr{X}_{s}$,
@@
    $\deg_{{\mathcal{L}}_{s}}(X^{\prime})<c(m,r,d,\deg_{{\mathcal{L}}_{s}^{\boxtimes m}}(\mathscr{X}_{S^{\prime}}^{[m]}\setminus\mathscr{X}_{S^{\prime}}^{[m],*})_{s})$, and
@@
The degree $\deg_{{\mathcal{L}}_{s}^{\boxtimes m}}(\mathscr{X}_{S^{\prime}}^{[m]}\setminus\mathscr{X}_{S^{\prime}}^{[m],*})_{s}$ is bounded from above solely in terms of $S^{\prime}$ and $m$. Hence (ii) implies a bound of the form $\deg_{{\mathcal{L}}_{s}}(X^{\prime})<c_{S^{\prime},2}^{\prime}$ with $c_{S^{\prime},2}^{\prime}$ depending only on $m$, $r$, $d$ and $S^{\prime}$. In summary, the set $\Sigma_{S^{\prime},s}$ defined in (4.7) is contained in $X^{\prime}(\overline{\mathbb{Q}})$ for a proper Zariski closed $X^{\prime}\subsetneq\mathscr{X}_{s}$ with $\deg_{{\mathcal{L}}_{s}}(X^{\prime})<c_{S^{\prime},2}^{\prime}$.
Letting $S^{\prime}$ runs over all irreducible components of $\overline{S_{\mathrm{gen}}}$, we set $c_{1}^{*\prime}:=\min_{S^{\prime}}\{c_{S^{\prime},1}^{\prime}\}>0$, $c_{2}^{*\prime}:=\max_{S^{\prime}}\{c_{S^{\prime},2}^{\prime}\}>0$ and $c_{3}^{*\prime}:=\max_{S^{\prime}}\{c_{S^{\prime},3}^{\prime}\}>0$. As the Zariski closed subset $\overline{S_{\mathrm{gen}}}$ of $S$ is determined by $S$, so are its irreducible components $S^{\prime}$. Moreover, $m=m(S)>0$ depends only on $S$. Therefore, the constants $c_{1}^{*\prime}$, $c_{2}^{*\prime}$ and $c_{3}^{*\prime}$ can be expressed only in terms of $r$, $d$ and $S$.
Let us summarize the result of the above discussion: For each $s\in\bigcup_{S^{\prime}}\pi^{[m]}(\mathscr{X}_{S^{\prime}}^{[m],*})(\overline{\mathbb{Q}})$, the set
@@
must be contained in $\Sigma_{S^{\prime},s}$ for some $S^{\prime}$; thus $\Sigma_{s}^{*}$ is contained in $X^{\prime}(\overline{\mathbb{Q}})$ for some proper Zariski closed $X^{\prime}\subsetneq\mathscr{X}_{s}$ with $\deg_{{\mathcal{L}}_{s}}(X^{\prime})<c_{2}^{*\prime}$. This concludes the proof in Case (ii). ∎
@@
Let $k$ be a field, $X$ an irreducible projective variety over $k$ and $L$ a very ample line bundle on $X$. Consider a closed subvariety $Z\subsetneq X^{M}$. Then there exists a constant
@@
such that, for any subset $\Sigma\subseteq X(k)$ satisfying $\Sigma^{M}\subseteq Z(k)$, there exists a proper Zariski closed $X^{\prime}\subsetneq X$ with $\Sigma\subseteq X^{\prime}(k)$ and $\deg_{L}(X^{\prime})<c$.
@@
We prove this lemma by induction on $M$. The base case $M=1$ is trivial as one may take $X^{\prime}:=Z$.
Assume next that the lemma is proved for $1,\ldots,M-1\geq 1$. Let $q\colon X^{M}\rightarrow X$ be the projection to the first factor. We write $Z^{\prime}$ for the union of all irreducible components $Y$ of $Z$ such that $q(Y)=X$ and $Z^{\prime\prime}$ for the union of the other components. Setting $\Sigma^{\prime\prime}:=q\left(\Sigma^{M}\cap Z^{\prime\prime}(k)\right)$, we notice that $(\Sigma\setminus\Sigma^{\prime\prime})\times\Sigma^{M-1}\subseteq Z^{\prime}(k)$. By [Gro67, Thm. 13.1.3 and Prop. 13.2.3], the fiber $q|_{Z^{\prime}}^{-1}(P)$ over a generic point $P\in X$ has dimension
$\dim Z^{\prime}-\dim X<M\cdot\dim X-\dim X=(M-1)\dim X,$
and the set
$W:=\{P\in X:\{P\}\times X^{M-1}\subseteq Z^{\prime}\}\subsetneq X$
is Zariski closed in $X$.
Case (i): $\Sigma\setminus\Sigma^{\prime\prime}\not\subseteq W(k)$.
Take a point $P\in\Sigma\setminus\Sigma^{\prime\prime}$ such that $P\notin W(k)$ and set $Z_{1}:=Z^{\prime}\cap(\{P\}\times X^{M-1})$. Since $\{P\}\times\Sigma^{M-1}\subseteq Z^{\prime}(k)$, we can apply the induction hypothesis for M1M-1, $\Sigma\subseteq X(k)$ and $Z_{1}\subseteq X^{M-1}$, tacitly identifying $\{P\}\times X^{M-1}$ with $X^{M-1}$. It implies the existence of a closed subvariety $X^{\prime}\subsetneq X$ such that $\Sigma\subseteq X^{\prime}(k)$ and
We claim that the constant on the right hand side can be expressed solely in terms of $M$, $\dim X$, $\deg_{L}(X)$ and $\deg_{L^{\boxtimes M}}(Z)$; this evidently suffices to complete the proof in this case. Indeed, we have
$\deg_{L^{\boxtimes(M-1)}}(Z_{1})=\deg_{L^{\boxtimes M}}(Z^{\prime}\cap(\{P\}\times X^{M-1}))$
and
by Bézout’s theorem. Using (2.6) inductively by taking $Y:=X$ and $Y^{\prime}:=X^{i}$ for i=1,2,,M2i=1,2,...,M-2, we have
$\deg_{L^{\boxtimes(M-1)}}(X^{M-1})=\frac{((M-1)\dim X)!}{((\dim X)!)^{M-1}}(\deg_{L}X)^{M-1}.$
Moreover $\deg_{L^{\boxtimes M}}(Z^{\prime})\leq\deg_{L^{\boxtimes M}}(Z)$. In conclusion, we get
Hence the right hand side of (4.9) depends only on $M$, $\dim X$, $\deg_{L}X$ and $\deg_{L^{\boxtimes M}}(Z)$. This proves the claim and ends the proof.
Case (ii): $\Sigma\setminus\Sigma^{\prime\prime}\subseteq W(k)$.
In this case, $(\Sigma\setminus\Sigma^{\prime\prime})\times X^{M-1}\subseteq Z^{\prime}$. We also assume that $k$ is algebraically closed, which we may by enlarging the base field if needed. As $Z^{\prime}\subsetneq X^{M}$, there exists a point $\underline{x}\in X^{M-1}(k)$ such that $(X\times\{\underline{x}\})\cap Z^{\prime}\not=X\times\{\underline{x}\}$. Thus $(X\times\{\underline{x}\})\cap Z^{\prime}=X^{\prime\prime}\times\{\underline{x}\}$ for some closed subset $X^{\prime\prime}\subsetneq X$. We have $\Sigma\setminus\Sigma^{\prime\prime}\subseteq X^{\prime\prime}(k)$ by construction, and
$\deg_{L}X^{\prime\prime}\leq\deg_{L^{\boxtimes M}}(X\times\{\underline{x}\})\deg_{L^{\boxtimes M}}(Z)=\deg_{L}X\deg_{L^{\boxtimes M}}(Z)$
by Bézout’s theorem.
To control the points in $\Sigma^{\prime\prime}$, note that the closed subset $q(Z^{\prime\prime})$ is properly contained in $X$ by definition. Moreover, we have
$\deg_{L}q(Z^{\prime\prime})\leq\deg_{L^{\boxtimes M}}(Z^{\prime\prime})\leq\deg_{L^{\boxtimes M}}(Z);$
the first inequality follows again from [Dil22, Lem. 2.4]. In summary, it suffices to take $X^{\prime}:=X^{\prime\prime}\cup q(Z^{\prime\prime})\subsetneq X$ in this case. ∎
4.2. The case of principally polarized abelian varieties
We next prove Proposition 4.1 in the special case of principally polarized abelian varieties.
Proposition 4.4.
Let $g$, $r$ and $d$ be positive integers. There exist constants $c_{1}^{\prime}=c_{1}^{\prime}(g,r,d)>0$, $c_{2}^{\prime}=c_{2}^{\prime}(g,r,d)>0$ and $c_{3}^{\prime}=c_{3}^{\prime}(g,r,d)>0$ satisfying the following property: For
  • every principally polarized abelian variety $(A,L)$ defined over $\overline{\mathbb{Q}}$ of dimension $g$,
  • and every irreducible subvariety $X$ of $A$ with $\dim X=r$ and $\deg_{L}X=d$ generating $A$,
the set
$\Sigma:=\left\{x\in X^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{L\otimes[-1]^{*}L}(x)\leq c_{1}^{\prime}\max\{1,h_{\mathrm{Fal}}(A)\}-c_{3}^{\prime}\right\}$
is contained in $X^{\prime}(\overline{\mathbb{Q}})$ for some proper Zariski closed $X^{\prime}\subsetneq X$ with $\deg_{L}(X^{\prime})<c_{2}^{\prime}$.
Proof.
For simplicity denote by $L_{-}:=[-1]^{*}L$.
@@ 4.2. The case of principally polarized abelian varieties
Let $X$ be an irreducible subvariety of $A$ generating $A$ such that $d=\deg_{L}X$ and $r=\dim X$. We may assume $r\geq 1$ because otherwise the result is trivial. We have
Consider the restricted Hilbert scheme $\mathbf{H}^{\circ}:=\mathbf{H}^{\circ}_{r,8^{r}d}(\mathfrak{A}_{g}/\mathbb{A}_{g})$ from (3.6) and recall the commutative diagram (4.2)
@@
of varieties and morphisms defined over $\overline{\mathbb{Q}}$. Note that $\pi|_{\mathscr{X}_{\mathbf{H}^{\circ}}}\colon\mathscr{X}_{\mathbf{H}^{\circ}}\rightarrow\mathbf{H}^{\circ}$ is the universal family.
The pair $(X,(A,L))$ is parametrized by a point $s\in\mathbf{H}^{\circ}(\overline{\mathbb{Q}})$ such that $X=\mathscr{X}_{s}$, $A=\mathcal{A}_{s}=(\mathfrak{A}_{g})_{b}$ and $\iota_{\mathbf{H}^{\circ}}(s)=b$.
@@
For the line bundle ${\mathcal{M}}=\overline{\iota_{\mathbf{H}^{\circ}}}^{*}\mathfrak{M}$ on $\overline{\mathbf{H}^{\circ}}$ defined above Proposition 4.2, we have
$h_{\overline{\mathbf{H}^{\circ}},{\mathcal{M}}}(s)=h_{\overline{\mathbb{A}_{g}},\mathfrak{M}}(\iota_{\mathbf{H}^{\circ}}(s))=h_{\overline{\mathbb{A}_{g}},\mathfrak{M}}(b).$
By fundamental work of Faltings [MB85, Thm. 1.1], $h_{\mathrm{Fal}}(A)$ is bounded from above in terms of $h_{\overline{\mathbb{A}_{g}},\mathfrak{M}}(b)=h_{\overline{\mathbf{H}^{\circ}},{\mathcal{M}}}(s)$ and $g$ only. More precisely,
(4.11) $\max\{1,h_{\mathrm{Fal}}(A))\}$ $\leq c(g)h_{\overline{\mathbf{H}^{\circ}},{\mathcal{M}}}(s)+c^{\prime}(g)\log(h_{\overline{\mathbf{H}^{\circ}},{\mathcal{M}}}(s)+2)$
$\leq(c(g)+c^{\prime}(g))h_{\overline{\mathbf{H}^{\circ}},{\mathcal{M}}}(s)+2c^{\prime}(g)$
for some constants $c(g)>0$ and $c^{\prime}(g)>0$.
@@
For the line bundle ${\mathcal{L}}=\iota^{*}\mathfrak{L}_{g}^{\otimes 4}$ on $\mathcal{A}_{\mathbf{H}^{\circ}}$ defined above Proposition 4.2, we have ${\mathcal{L}}|_{\mathcal{A}_{s}}=\mathfrak{L}_{g}^{\otimes 4}|_{(\mathfrak{A}_{g})_{\iota(s)}}=\mathfrak{L}_{g}^{\otimes 4}|_{(\mathfrak{A}_{g})_{b}}$, and hence ${\mathcal{L}}|_{A}=\mathfrak{L}_{g}^{\otimes 4}|_{A}$ via the identification $A=\mathcal{A}_{s}=(\mathfrak{A}_{g})_{b}$ fixed above. Thus $c_{1}({\mathcal{L}}|_{A})=4c_{1}(L\otimes L_{-})$. Moreover, both ${\mathcal{L}}|_{A}$ and $(L\otimes L_{-})^{\otimes 4}$ are symmetric and ample on $A$, so Lemma 2.3 implies